OpenStudy (pgxgirl103):

Solve for x? √ (5x-1) = √ (4x+9)

OpenStudy (anonymous):

square both sides (just like before) .

OpenStudy (pgxgirl103):

do i have to do it separately? since there are 2.

OpenStudy (anonymous):

Well, if you had √(x+1) = √(2x-2) after squaring both sides, the square roots would just disappear, and you would get x+1=2x-2 then.. 1=x-2 x=3 So, what happens if you square both sides, in YOUR problem ?

OpenStudy (pgxgirl103):

I have no idea. I'm confused because there are 2.

OpenStudy (anonymous):

raise each side to the second power.... lets start with the left side √ (5x-1) = √ (5x-1) ² (raised to 2nd power) = 5x-1. now continue with the right side √ (4x+9) = √ (4x+9) ² (raised to 2nd power) = 4x+9. So, after raising BOTH SIDES to the SECOND POWER, you would get 5x-1=4x+9 See? Can you take it from here ?

OpenStudy (pgxgirl103):

do you just subtract?

OpenStudy (anonymous):

After squaring both sides? Yes... 1) subtract 4x from both sides 2) add 1 to both sides. The answer is ?

OpenStudy (pgxgirl103):

5-4 is 1 and 4-4 is 0. do you put those together to get 10?

OpenStudy (anonymous):

I don't really get what you said by 4-4 =0, but yeah, the ans. is 10 .

OpenStudy (pgxgirl103):

well how do you understand it then?

OpenStudy (pgxgirl103):

I don't want to do something wrong and mess up on the others

OpenStudy (anonymous):

5x-1=4x+9 -4x -4x x-1=9 +1 +1 x=10